OpenStudy (anonymous):

What is the equation of a line that passes through (7, -2) and has a slope of 3?

OpenStudy (anonymous):

yo

OpenStudy (anonymous):

Use point-slope form. \[y - y_1 = m(x - x_1)\] m = 3 (x1, y1) = (7, -2)

OpenStudy (amistre64):

which "equation" of the line is appropriate?

OpenStudy (anonymous):

first plug in the slope to your point-slope equation to get: y=3x+b now put in your point for x and y and solve for b -2=21+b b=-23 so your equation is: y=3x-23

OpenStudy (anonymous):

i mean slope-intercept

OpenStudy (amistre64):

I like the: y = mx -mPx +Py version :)

OpenStudy (amistre64):

y = 3x -3(7) -2